Find an equation of a line that is perpendicular to the line of 5x-4y=20.
I had to find the x and y intercepts. I have x=5 and y=4. My answer needs to be in y=mx+b form.

Found 2 solutions by stanbon, sister_insomnia:
Answer by stanbon(75887)

Find an equation of a line that is perpendicular to the line of 5x-4y=20.
I had to find the x and y intercepts. I have x=5 and y=4. My answer needs to be in y=mx+b form.
----------
Find the slope of the given line, as follows:
5x-4y=20
4y=5x-20
y=(5/4)x-5
The given line has a slope of 5/4
So the line perpendicular to it must have a slope of -4/5
Therefore the equation of a line perpendicular to the
given line would be y=(-4/5)x+b where b is any number you
wish to make it.
Cheers,
Stan H.

Answer by sister_insomnia(8)

5x-4y=20
Start by putting this equation into y=mx+b:
5x-4y=20
Subtract 5x:
-4y=-5x+20
Divide by -4:
y=5/4x-5
Now that your equation is in y=mx+b, you can easily identify the slope, or m. Since m is 5/4, that is your slope. Remember the slope of perpendicular lines is the negative reciprocal. Therefore, the slope of the perpendicular line is-4/5. y=-4/5-5 is the equation of the perpendicular line.